Question 8 The nylon rope from which a mountain climber is suspended has a force constant of 1.40 x 104 N/m. How much would this rope stretch to break the climbers fall if he free-falls 2.16 m before the rope runs out of slack, given his mass plus and the mass of his equipment are 91.6 kg? Hint: Use conservation of energy. You can assume the amount the rope stretches is small compared to the distance of his fall. Select the correct answer O 52.6cm O 46.2cm O 58.7 cmAnswer O 43.5 cm O 46.0 cm CHECK ANSWER of 2 attempts used LAST ATTEMPT!

Question9 A 0.151-kg ball is hung vertically from a spring. The spring stretches by 2.27 cm from its natural length when the ball is hanging at equilibrium. A child comes along and pulls the ball down an additional 5cm, then lets go. How long (in seconds) will it take the ball to swing up and down exactly 4 times, making 4 complete oscillations before again hitting its lowest position? Hints: This question contains more information than necessary. Also, beware of unit conversions! Enter answer here 0.25 seconds 0.25secondsXXSwr 1 of 3 attempts used CHECK ANSWER

Question 10 This question contains multiple parts. Make sure to read all the instructions and answer each part Part a What will happen to the period of a ball-on-a-spring if the ball is replaced by a smaller one with half the mass? Select the correct answer O Period stays the same O Period is halved Period decreases by 29% Period is doubled Your Answer Period increases by 41% CHECK ANSWER 1 of 2 attempts used LAST ATTEMPT
Answer #1

As HOMEWORKLIB RULES guide lines i did 8 only

8)

from conservation of energy

PEelastic =PEgravitational

1/2*k*x^2 =m*g*h

k*x^2 =2*m*g*h

x^2 =2*m*g*h/(k)

x = sqrt((2*m*g*h)/(k))

=sqrt((2*91.6 kg*9.80 m/s^2*2.16 m)/(1.40*10^4 N/m))

x = 0.52630 m

x = 52.6 cm

option A is correct